OpenStudy (anonymous):

What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374? A. 11 B. 11.5 C. 12 D 12.5

OpenStudy (anonymous):

How to do this wasn't in my text book

OpenStudy (anonymous):

here n=43 first term a=-13 sum s=9374 s=43/2[2(-13)+(43-1)d] 9374=43/2[-26+42d] 9374*2/43=-26+42d 18748/43=-26+42d 436+26=42d 42d=462 d=462/42=11

OpenStudy (anonymous):

So the answer is 11?

OpenStudy (anonymous):

can you help me with anthor one?